OpenStudy (anonymous):

For what value of x is sin x = cos 19°, where 0°< x < 90°?

OpenStudy (anonymous):

Since u have asked sin x=1/2 According to identity, Sin2 x +cos2 x =1 Cos 2 x =1 − sin 2x Cos x = √1−sin 2x Cos x =√1−(1/2)2 Cos x =√1−1/4 Cos x =√3/4 Cos x =√3/2 By this way u can find the cos of given sin and vice versa because because it is not possible that we always know the value of given trigonometric function orally eg. Sin x = 1/10 etc……. Hope u understand what i want to say

OpenStudy (anonymous):

TA DA

OpenStudy (anonymous):

!!!!!

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

no i dont

OpenStudy (jdoe0001):

hint: \(\bf \textit{Cofunction Identities} \\ \quad \\ sin\left(\frac{\pi}{2}-{\color{red}{ \theta}}\right)=cos({\color{red}{ \theta}})\qquad cos\left(\frac{\pi}{2}-{\color{red}{ \theta}}\right)=sin({\color{red}{ \theta}})\) any ideas?

OpenStudy (jdoe0001):

or lemme put it differently \(\bf sin\left(90^o-{\color{red}{ \theta}}\right)=cos({\color{red}{ \theta}})\qquad cos\left(90^o-{\color{red}{ \theta}}\right)=sin({\color{red}{ \theta}}) \)